# Enhancing High-Level Synthesis with Automated Pragma Insertion and Code Transformation Framework

### v3

@@ Abstract.
High-level synthesis, source-to-source compilers, and various Design Space Exploration techniques for pragma insertion have significantly improved the Quality of Results of generated designs. These tools offer benefits such as reduced development time and enhanced performance. However, achieving high-quality results often requires additional manual code transformations and tiling selections, which are typically performed separately or as pre-processing steps. Although DSE techniques enable code transformation upfront, the vastness of the search space often limits the exploration of all possible code transformations, making it challenging to determine which transformations are necessary. Additionally, ensuring correctness remains challenging, especially for complex transformations and optimizations.
To tackle this obstacle, we first propose a comprehensive framework leveraging HLS compilers. Our system streamlines code transformation, pragma insertion, and tiles size selection for on-chip data caching through a unified optimization problem, aiming to enhance parallelization, particularly beneficial for computation-bound kernels. Them employing a novel Non-Linear Programming (NLP) approach, we simultaneously ascertain transformations, pragmas, and tile sizes, focusing on regular loop-based kernels. Our evaluation demonstrates that our framework adeptly identifies the appropriate transformations, including scenarios where no transformation is necessary, and inserts pragmas to achieve a favorable Quality of Results.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 27–31, 2024; New Jersey, USAisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
High-level synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code. Despite the improvements in productivity, both code transformation and ensuring the correctness of the transformations remains challenging. The HLS design space encompasses various code transformations and pragma placements. However, exploring this joint space can be daunting due to its vastness, with millions of potential points, and traditional analytical methods struggle to efficiently navigate it. Key issues include the lack of convexity and regularity in this space. Moreover, separating code transformation and pragma insertion into distinct optimization problems complicates the search and may necessitate initiating the search for one optimization problem upon any changes occurring in the other.
@@
Our primary research objective is to develop a system capable of autonomously conducting search-friendly code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization problem. Even though several exploration methods could be used, we chose to employ a Non-Linear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We develop an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
@@
  • A novel template that streamlines code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • We develop a novel NLP approach tailored to exploring this joint space, particularly focusing on regular loop-based kernels. The NLP allows for the exploration of a space whose resources are set by the user.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE and HARP, all without the necessity of Design Space Exploration. Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
The paper is organized into the following sections. Section 2 lays out the motivation behind our approach and the proposed solution. Following that, Section 3 presents the range of code transformation and pragma insertion we consider. In Section 4, we introduce a non-linear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Moving forward, Section 5 elaborates on our code generation process. Finally, sections 6, 7 and 8 are dedicated to evaluating our method, presenting related work, and drawing conclusions.
@@ 2.1. Space to explore
The exploration space for optimizing an HLS design encompasses pragma insertion, code transformation, and data caching, all while adhering to resource constraints. While various objective functions can be considered, our focus in this work is on minimizing latency by maximizing parallelism. This objective function is especially potent for computation-bound kernels, addressing the bottleneck caused by insufficient parallelization. These different elements of the space (pragma insertion, code transformation, and data caching) are inherently interconnected. Inserting unrolled pragmas influences array partitioning and thus the utilization of BRAM. The schedule and pragmas influence the possibility of caching data on-chip, and the choice of schedule impacts how we want to unroll based on loop properties (e.g., reduction loop), and so on. Thus, the choice of one element affects the others, necessitating backtracking if the spaces are separated. This complexity complicates the exploration of the space and also requires constant verification of the legality of transformations.
@@ 2.2. Limitation of the current DSE
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022) enable code transformations and pragma insertion, but their spaces are limited for both. Transformations are restricted to loop property-based permutations or the use of Pluto (Bondhugula et al., 2008), which tiles the program but may restrict parallelization for FPGAs and does not consider FPGA-specific optimizations such as pipelining or array partitioning.
On the other hand, numerous Design Space Explorations (DSEs) (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragma insertion for a fixed loop order. Although the user may initially set the schedule, it can be difficult, even for experts, to predict which transformations are necessary prior to pragma insertion. Furthermore, the multitude of code transformations renders the exploration of each pragma impractical in terms of scalability.
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a). A comparison with HARP (Sohrabizadeh et al., 2023) will be presented in Section 6. AutoDSE trents the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD/Xilinx source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones.
In the upcoming examples, we utilize the HLS compiler Vitis 2023.2, excluding the unsafe-math option. We examine two scenarios: the kernel gemm from Polybench (Polybench, [n. d.]) as shown in Listing 1, and CNN.
1for (i = 0; i < 200; i++) {
2 for (j = 0; j < 220; j++)
3 C[i][j] *= beta;
4 for (k = 0; k < 240; k++)
5 for (j = 0; j < 220; j++)
6 C[i][j] += alpha * A[i][k] * B[k][j];}
Listing 1: Code original of gemm
The gemm kernel’s original loop order, as illustrated in Listing 1, situates a non-reduction loop innermost for the second statement. This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop. However, partially unrolling the reduction loop in this scenario diminishes performance due to the increased initiation interval (II) of the pipeline resulting from unrolling the reduction loop. As a consequence, the designs yielding the best QoR for these two methods achieve a throughput of approximately 20 GF/s. Listing 1 highlights the design found by AutoDSE. The pragmas with $factor=1$ or pipeline off are not utilized in the final design. However, they illustrate the potential space explored by AutoDSE with various factors. In such instances, a transformation becomes imperative to augment parallelization.
1#pragma ACCEL parallel factor=1
2#pragma ACCEL pipeline off
3#pragma ACCEL TILE FACTOR=5
4for (i = 0; i < 200; i++) {
5#pragma ACCEL tile factor=1
6#pragma ACCEL pipeline off
7#pragma ACCEL parallel FACTOR=4
8 for (j = 0; j < 220; j++)
9 C[i][j] *= beta;
10#pragma ACCEL tile factor=1
11#pragma ACCEL parallel factor=1
12#pragma ACCEL pipeline flatten
13 for (k = 0; k < 240; k++)
14#pragma ACCEL pipeline off
15#pragma ACCEL tile factor=1
16 for (j = 0; j < 220; j++) // fully unrolled
17 C[i][j] += alpha * A[i][k] * B[k][j];}
Listing 2: Code of gemm found by AutoDSE
When dealing with Convolutional Neural Networks (CNNs), there’s significant room for code optimization, especially considering the need for tiling. Each loop split introduces various permutations, leading to a vast range of options, totaling $2.23\times 10^{13}$ potential combinations of loop orders and tile sizes. Despite achieving satisfactory QoR — 42.15 GF/s for AutoDSE and 31.80 GF/s for NLP-DSE — both methods fall short of fully exploiting the available level of parallelism, hinting at further optimization potential. However, exhaustively exploring each code transformation proves practically infeasible due to the sheer volume of possibilities. While the Merlin compiler can leverage strip mining to enhance on-chip data caching potential, its capabilities in this regard are constrained by the inability to permute the loops.

### v4

@@ Abstract.
High-Level Synthesis compilers and Design Space Exploration tools have greatly advanced the automation of hardware design, improving development time and performance. However, achieving a good Quality of Results still requires extensive manual code transformations, pragma insertion, and tile size selection, which are typically handled separately. The design space is too large to be fully explored by this fragmented approach. It is too difficult to navigate this way, limits the exploration of potential optimizations, and complicates the design generation process.
To tackle this obstacle we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a single optimization problem. By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
Evaluation against state-of-the-art frameworks, including AutoDSE, NLP-DSE, and ScaleHLS, shows that Sisyphus achieves superior Quality of Results, outperforming alternatives across multiple benchmarks. By integrating code transformation and pragma insertion into a unified model, Sisyphus significantly reduces design generation complexity and improves performance for FPGA-based systems.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; March xx–xx, 2025; California, USAisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024a, b; Sohrabizadeh et al., 2021, 2023, 2022; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code.
Despite their advantages, existing HLS workflows face several significant challenges. Previous frameworks (Ye et al., 2022, 2024; Lai et al., 2019; Zhang et al., 2024) have integrated pragma insertion and code transformation, but these efforts are somewhat limited, focusing primarily on loop permutations based on loop properties. Additionally, code transformations and pragma insertion are often treated as separate processes. Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, while pragmas for optimizations like loop unrolling, pipelining, and array partitioning are applied later. This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa. Moreover, the immense design space—encompassing millions of potential transformation and pragma combinations—makes exhaustive exploration unfeasible. Ensuring the correctness of these transformations while adhering to resource constraints further complicates the process.
@@
Our primary research objective is to develop a system capable of autonomously conducting code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization framework. Even though several exploration methods could be used, we chose to employ a Nonlinear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2024b, a). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
@@
  • A novel template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels. Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within seconds or minutes. Additionally, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP, all without the necessity of DSE. Furthermore, Sisyphus accurately identifies the appropriate transformation, even in cases where no transformation is required.
The paper is organized into the following sections. Section 2 lays out the motivation behind our approach and the proposed solution. Following that, Section 3 presents the range of code transformation and pragma insertion we consider. In Section 4, we introduce a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Section 5 elaborates on our post-optimizations. Finally, sections 6, 7, 8 and 9 are dedicated to evaluating our method, presenting related work, addressing limitations and drawing conclusions.
@@ 2.1. Space to Explore
The exploration space for optimizing an HLS design includes pragma insertion, code transformations, and data caching, all within the limits of resource constraints. Although different objective functions can be explored, this work specifically focuses on minimizing latency while respecting resource constraints. This objective function is especially potent for computation bound kernels because it addresses the bottleneck caused by insufficient parallelization.
The various components of the exploration space—pragma insertion, code transformations, and data caching—are inherently interdependent. For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization. Similarly, the scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops. A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1). The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
Consequently, selecting one element impacts the others, often requiring backtracking when the spaces are treated separately. This interconnection adds complexity to the exploration process and necessitates ongoing verification to ensure the legality of transformations.
1for (i = 0; i < 200; i++) { // L0
2 for (j = 0; j < 220; j++) // L1
3 C[i][j] *= beta; // S0
4 for (k = 0; k < 240; k++) // L2
5 for (j = 0; j < 220; j++)// L3
[…]
9for (i = 0; i < 200; i++) // L0
10 for (j = 0; j < 220; j++) // L1
11 C[i][j] *= beta; // S0
12for (i = 0; i < 200; i++) // L2
13 for (k = 0; k < 240; k++) // L3
14 for (j = 0; j < 220; j++) // L4
[…]
18for (i0 = 0; i0 < I0_S0; i0++) // L0
19 for (i1 = 0; i1 < I1_S0; i1++) // L1
20 for (i2 = 0; i2 < I2_S0; i2++) // L2
21 for (j0 = 0; j0 < J0_S0; j0++) // L3
22 for (j1 = 0; j1 < J1_S0; j1++) // L4
23 for (j2 = 0; j2 < J2_S0; j2++) // L5
24 C[i][j] *= beta; // S0
25for (i = 0; i < I0_S1; i++) // L6
26 for (i = 0; i < I1_S1; i++) // L7
27 for (i = 0; i < I2_S1; i++) // L8
28 for (k = 0; k < K0_S1; k++) // L9
29 for (k = 0; k < K1_S1; k++) // L10
30 for (k = 0; k < K2_S1; k++) // L11
31 for (j = 0; j < J0_S1; j++) // L12
32 for (j = 0; j < J1_S1; j++) // L13
33 for (j = 0; j < J2_S1; j++) // L14
[…]
38
[…]
42On-chip mem = 320 kB}]
43load_C(C, C_off_chip);
44/**** Level 0 of S0 ****/
45for (j0 = 0; j0 < 1; j0++) // L0
46 for (i0 = 0; i0 < 4; i0++) // L1
47/**** Level 1 of S0 ****/
48 for (j1 = 0; j1 < 22; j1++) // L2
49#pragma HLS pipeline
50/**** Level 2 of S0 ****/
51 for (j2=0; j2<10; j2++) { // L3
52#pragma HLS unroll
53 for (i2=0; i2<50; i2++) // L4
54#pragma HLS unroll
55 i = ...;
56 C[i][j] *=beta; } // S0
57/**** Level 0 of S1 ****/
58for (k0 = 0; k0 < 48; k0++) // L5
59 load_B_S1(B, B_off_chip, k0);
60 load_A_S1(A, A_off_chip, k0);
61 for (j0 = 0; j0 < 1; j0++) // L6
62 for (i0 = 0; i0 < 4; i0++) // L7
63/**** Level 1 of S1 ****/
64 for (j1=0; j1 < 220; j1++) // L8
65#pragma HLS pipeline
66/**** Level 2 of S1 ****/
67 for (j2=0; j2<1; j2++) // L9
68#pragma HLS unroll
69 for (i2=0; i2<50; i2++) // L10
70#pragma HLS unroll
71 for(k2=0;k2<5;k2++){ // L11
72#pragma HLS unroll
73 i = ...;
74 C[i][j]+=alpha*A[i][k2]*B[k2][j];}// S1
75store_C(C, C_off_chip);
Listing 1: Original Code
1load_C(C, C_off_chip);
2load_A(A, a_off_chip);
3load_B(B, B_off_chip);
4/**** Level 0 of S0 ****/
5for (i0 = 0; i0 < 1; i0++) // L0
6 for (j0 = 0; j0 < 1; j0++) // L1
7/**** Level 1 of S0 ****/
8 for (j1 = 0; j1 < 55; j1++) // L2
9#pragma HLS pipeline
10/**** Level 2 of S0 ****/
11 for (i2 = 0; i2 < 200; i2++) // L3
12#pragma HLS unroll
13 for (j2=0; j2<4; j2++) { // L4
14#pragma HLS unroll
15 i = ...;
16 C[i][j] *= beta;} // S0
17/**** Level 0 of S1 ****/
18for (i0 = 0; i0 < 1; i0++) // L5
19 for (j0 = 0; j0 < 1; j0++) // L6
20 for (k0 = 0; k0 < 60; k0++) // L7
21/**** Level 1 of S1 ****/
22 for (j1 = 0; j1 < 220; j1++) // L8
23#pragma HLS pipeline
24/**** Level 2 of S1 ****/
25 for (i2=0;i2<200;i2++) // L9
26#pragma HLS unroll
27 for (j2=0; j2<1; j2++) // L10
28#pragma HLS unroll
29 for(k2=0;k2<4;k2++){ // L11
30#pragma HLS unroll
31 i = ...;
32 C[i][j]+=alpha*A[i][k]*B[k][j];}//S1
33store_C(C, C_off_chip);
Listing 2: Result found using the NLP with the constraints:
DSP = 6,840 On-chip mem = 7.2 MB
@@ 2.2. Limitation of the Current DSE
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024). Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning. Additionally, Pluto minimizes memory transfer from off-chip to on-chip (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat). Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2021, 2023, 2022; Pouget et al., 2024b) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging. Exploring every possible transformation is also impractical for large design spaces. The key differences between the frameworks are summarized in Table 1.
Poly-Opt-HLS POL-SCA/ Pluto Auto-DSE/ HARP Hetero-CL / Allo NLP-DSE Scale-HLS/ POM Sisy-phus
Tiling Limit.
Permutation Limit. Limit. Limit.
Pragma Insert.
Code Trans + Pragma Insert. (unified)
Off-chip comm. gen.
Enum. (AI, heuristics,…) Manual
Objective Comm Comm Lat - Lat Lat Lat
Table 1. Comparison of different frameworks
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2021) and NLP-DSE (Pouget et al., 2024b, a). A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6. AutoDSE treats the compiler Merlin (Xilinx, 2024) as a black box, adjusting pragmas based on identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD/Xilinx source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones. A more detailed description of the Merlin compiler can be found in (Pouget et al., 2024b). In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled. The targeted FPGA is the AMD/Xilinx Alveo U200. We evaluate two scenarios: the gemm kernel from Polybench (Polybench, [n. d.]), as shown in Listing 1, and a convolutional neural network (CNN) layer where the problem size and original loop order of CNN are I,J=256 (Output and Input Channels), H,W=224 (Height, Weight) and P,Q=5 (Filter Height and Weight). The gemm kernel’s original loop order, as illustrated in Listing 1, has a non-reduction loop at the center of the second statement (L3). This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop (L2). In this scenario, if the reduction loop is pipelined and partially unrolled by a factor of $uf$, performance will degrade. While partial unrolling allows for parallel execution of the multiplications, it increases the pipeline depth due to the loop’s dependencies, as the $uf$ additions must still be performed sequentially. Consequently, when the reduction loop is pipelined, the initiation interval (II) will significantly increase, becoming a multiple of both $uf$ and the latency of the reduction operation. This occurs because the next iteration of the loop can only begin after the $uf$ additions are completed and the output is fully written. Their inability to further increase parallelism due to loop order limits the best designs to a throughput of 20 GFLOPs per second (GF/s). In such instances, a transformation becomes imperative to augment parallelization. A single CNN layer with 6 loops results in an enormous design space that cannot be explored exhaustively. Loop permutations alone represent 6! = 720 possibilities, and tiling is required to manage array sizes exceeding on-chip memory capacity. While frameworks like AutoDSE and NLP-DSE achieve satisfactory throughput—42.15 GF/s and 31.80 GF/s, respectively—they are unable to explore the full design space, which includes both code transformations and pragma insertion, within a reasonable timeframe. For example, running AutoDSE for all possible loop permutations would take around 500 days, and for NLP-DSE, it would take 106 days. Other frameworks, which apply code transformations based solely on loop properties, such as prioritizing reduction loops in the outermost positions, explore a much more restricted design space, limiting their potential for further optimization. This underscores the challenge of navigating such a vast design space and highlights gaps in current approaches.